# Diffusion-Free Dynamics in Rotating Spherical Shell Convection Driven By Internal Heating and Cooling

### v1

@@ 3.2.2 Reynolds number scaling
under the assumption that the velocity $U$ (used to compute $Re$) and the temperature difference $\Delta T$ (used to compute $Ra$) are independent of the diffusivities. When instead the system is rotating, then a diffusion-free balance between the Coriolis, inertial, and Archimedean (buoyancy driving) terms in the momentum equation (CIA balance) suggests the scaling (J. Aubert et al., 2001; A. J. Barker et al., 2014; T. Gastine et al., 2016):
@@
sometimes referred to as the inertial scaling of rotating convection. If alternatively, a balance is sought between the Coriolis, viscous, and Archimedean terms (i.e., VAC balance; E. M. King et al., 2013), then the following scaling:
@@
is obtained. Equation 17 implies that $U$ depends on the diffusivity. We are not aware of any study that recovers either of the diffusion-free scalings (Equations 15 and 16) from simulations of spherical shell convection that include rotation. For example, even the simulations presented in T. Gastine et al. (2016), that achieve an RMLT-like scaling for $Nu$ in the rapidly rotating regime, exhibit a Reynolds number scaling that is not diffusion-free. There the rapidly rotating cases follow an alternate scaling in which CIA balance is obtained in the bulk, but kinetic energy dissipation is dominated by the boundary layers, while their weakly-nonlinear simulations follow the VAC scaling described above. However, we note that N. A. Featherstone & B. W. Hindman (2016) do find diffusion-free behaviour for the kinetic energy in simulations that are non-rotating.
Figure 6c and 6d present a comparison between the Reynolds number scaling obtained in our simulations, and the theoretical relations given by Equations 1517. As with the Nusselt number, the three regimes we have identified correspond to three distinct scaling relations. Experiments with $Ro_{\text{cv,F}}>1$ in the rotationally-uninfluenced regime, for which small scales are rotationally-unconstrained, follow the ultimate scaling given by Equation 15 (green points in Figure 6d). Meanwhile, the experiments that are rotationally-influenced ($Ro_{\text{bulk}}<1$ but $Ro_{\omega}\gtrsim 1$) follow the inertial scaling (blue points in Figure 6d). By contrast, the rotationally-constrained $Ro_{\text{cv,F}}<0.1$ cases display a Reynolds number scaling that closely follows the diffusivity-dependent VAC balance scaling (orange points in Figure 6c), which is consistent with the dependence of the KE on the diffusivities in Figure 2.
@@ 4 Summary and Discussion
The scaling results presented in the previous section suggest that our simulations can be grouped into three distinct regimes, which we term rapidly rotating ($Ro_{\text{cv,F}}<0.1$), rotationally-influenced ($0.1\leq Ro_{\text{cv,F}}<1$), and rotationally-uninfluenced ($Ro_{\text{cv,F}}\geq 1$). In the rotationally-constrained regime, both the small scales and the bulk are constrained by rotation. The rotationally-influenced regime is an intermediate regime where large scales are constrained by rotation, but small scales are not (Figure 1d). Finally, there is no discernible impact of rotation on the dynamics of simulations in the rotationally-uninfluenced regime. The key features of each regime are summarised in Table 1 and described below.
@@
Previous work has shown that ‘local Nusselt numbers’ defined for the equatorial and polar regions can yield different scaling results that are set by the local dynamics (T. Gastine & J. M. Aurnou, 2023). Therefore, the fact that RMLT does not fully capture the convective heat transport scaling in the rapidly-rotating experiments is not surprising, given that the morphology of the convection depends strongly on latitude (e.g., wave-like dynamics at low latitudes but convective turbulence at high latitudes) but this effect is not accounted for by RMLT. A full analysis of the latitudinal heat budget, as well as the dependence of the size of the equatorial region on $Ro_{\text{cv,F}}$, is required to form a complete understanding of the $Nu$ scaling in the rapidly-rotating regime. A second, puzzling feature of the rapidly-rotating regime is that the convective velocity scaling (i.e., the $Re$ scaling) is diffusivity-dependent, but the heat transport scaling is not. Our analysis in this article has focused on the CZ in isolation; however, by analogy with the influence that boundary layers can have on the bulk in boundary-driven convection (S. Grossmann & D. Lohse, 2000; E. M. King et al., 2012; T. Gastine et al., 2016), it is plausible that the dynamics of the heated and cooled regions can exert an influence on the CZ. We note that forming a volume-integrated relation between $Re$ and the viscous dissipation rate is one route to obtaining the inertial scaling for $Re$ (Equation 15; T. Gastine et al., 2016). As we are interested in the dynamics integrated over the CZ, this integral relation will not be exact, but instead contain ‘boundary terms’ that describe the exchange of kinetic energy between the CZ and the heated and cooled regions. Understanding the details of this exchange are likely important to understanding the scaling behaviour of $Re$ in the CZ.
@@
Several aspects of the flows here are unlike those realised in a star like the Sun. We believe most of these discrepancies result from choices and simplifications made in our modeling (rather than from a direct dependence on diffusivity), but testing this assertion will require future work. For example, none of the zonal flows in our calculations are particularly ‘Solar-like’; those that exhibit a fast equator typically also have a ‘polar vortex’, consisting of retrograde flow at high latitudes. This may reflect the narrow aspect ratio of the convective shell considered here (see, e.g., P. A. Gilman, 1979). More generally, we have restricted ourselves to the simplest case of Boussinesq, hydrodynamic convection. This set-up has the advantage of being easily relatable to theory. However, both magnetism and compressibility have important effects on the dynamics of rotating, spherical shell convection (e.g., T. Gastine & J. Wicht, 2012; R. K. Yadav et al., 2016; P. J. Käpylä, 2023; H. Hotta et al., 2015, 2022; K. M. Soderlund et al., 2025). Therefore, incorporating these effects is an obvious and important follow-up to this work. In addition, we have not analysed the sensitivity of our results to the functional form of the prescribed internal heating and cooling profile. We intend to study these effects in future work.
@@
We thank Benjamin Brown and Keaton Burns for useful suggestions with respect to configuring and using Dedalus. NTL and TJ-H benefited from stimulating discussions at the workshop Rotating Turbulence: Interplay and Separability of Bulk and Boundary Dynamics, hosted by Institute for Pure and Applied Mathematics (IPAM) at the University of California, Los Angeles. IPAM is supported by the National Science Foundation (grant no. DMS-1925919). MKB thanks the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Frontiers in Dynamo Theory: From the Earth to the Stars (DYT2), where preliminary work on this topic was conducted; this was supported by EPSRC grant no EP/R014604/1. MKB also thanks Evan Anders, and other participants in the DYT2 programme, for useful discussions. NTL, TJ-H, MKB, LKC, and ST received funding from the STFC under grant agreements ST/Y002156/1, ST/Y002296/1, and ST/Y002113/1. LKC also gratefully acknowledges support from the STFC under grant agreement ST/X001083/1, and TJ-H was additionally supported by an STFC Studentship under grant agreement ST/W507453/1. This work used the DiRAC Memory Intensive service (Cosma8) at Durham University, managed by the Institute for Computational Cosmology, and the DiRAC Data Intensive service (DIaL3) at the University of Leicester managed by the University of Leicester Research Computing Service. These facilities are managed on behalf of the STFC DiRAC HPC (www.dirac.ac.uk). The DiRAC services at Durham and Leicester were funded by BEIS, UKRI and STFC capital funding, and STFC operations grants. The service at Durham received funding from Durham University. DiRAC is part of the UKRI Digital Research Infrastructure. {contribution}NTL, MKB, LKC, and ST designed the research. NTL performed the numerical simulations and data analysis. All authors contributed to the interpretation of results and to writing the manuscript.

### v2

@@ 3.2.2 Reynolds number scaling
under the assumption that the velocity $U$ (used to compute $Re$) and the temperature difference $\Delta T$ (used to compute $Ra$) are independent of the diffusivities. When instead the system is rotating, then a diffusion-free balance between the Coriolis, inertial, and Archimedean (buoyancy driving) terms in the momentum equation (CIA balance) suggests the scaling (2001PEPI..128...51A; 2014ApJ...791...13B; 2016JFM...808..690G; 2021PNAS..11822518V):
@@
sometimes referred to as the inertial scaling of rotating convection. If alternatively, a balance is sought between the Coriolis, viscous, and Archimedean terms (i.e., VAC balance; 2013JFM...717..449K), then the following scaling:
@@
is obtained. Equation 17 implies that $U$ depends on the diffusivity. We are not aware of any study that recovers either of the diffusion-free scalings (Equations 15 and 16) from simulations of spherical shell convection that include rotation. For example, even the simulations presented in 2016JFM...808..690G, that achieve an RMLT-like scaling for $Nu$ in the rapidly rotating regime, exhibit a Reynolds number scaling that is not diffusion-free. There the rapidly rotating cases follow an alternate scaling in which CIA balance is obtained in the bulk, but kinetic energy dissipation is dominated by the boundary layers, while their weakly-nonlinear simulations follow the VAC scaling described above. However, we note that 2016ApJ...818...32F do find diffusion-free behaviour for the kinetic energy in simulations that are non-rotating.
Figure 6c and 6d present a comparison between the Reynolds number scaling obtained in our simulations, and the theoretical relations given by Equations 1517. As with the Nusselt number, the three regimes we have identified correspond to three distinct scaling relations. Experiments with $Ro_{\text{cv,F}}>1$ in the rotationally-uninfluenced regime, for which all scales are rotationally-unconstrained, follow the ultimate scaling given by Equation 15 (green points in Figure 6d). Meanwhile, the experiments that are rotationally-influenced ($Ro_{\text{bulk}}<1$ but $Ro_{\omega}\gtrsim 1$) follow the inertial scaling (blue points in Figure 6d). By contrast, the rotationally-constrained $Ro_{\text{cv,F}}<0.1$ cases display a Reynolds number scaling that closely follows the diffusivity-dependent VAC balance scaling (orange points in Figure 6c), which is consistent with the dependence of the KE on the diffusivities in Figure 2.
@@ 4 Summary and Discussion
The scaling results presented in the previous section suggest that our simulations can be grouped into three distinct regimes, which we term rotationally-constrained ($Ro_{\text{cv,F}}<0.1$), rotationally-influenced ($0.1\leq Ro_{\text{cv,F}}<1$), and rotationally-uninfluenced ($Ro_{\text{cv,F}}\geq 1$). In the rotationally-constrained regime, all spatial scales are constrained by rotation. The rotationally-influenced regime is an intermediate regime where largest scales are constrained by rotation, but some intermediate scales are not (Figure 1d). Finally, there is no discernible impact of rotation on the dynamics of simulations in the rotationally-uninfluenced regime. The key features of each regime are summarised in Table 1 and described below.
@@
Previous work has shown that ‘local Nusselt numbers’ defined for the equatorial and polar regions can yield different scaling results that are set by the local dynamics (2023JFM...954R...1G). Therefore, the fact that RMLT does not fully capture the convective heat transport scaling in the rotationally-constrained experiments is not surprising, given that the morphology of the convection depends strongly on latitude (e.g., wave-like dynamics at low latitudes but convective turbulence at high latitudes) but this effect is not accounted for by RMLT. A full analysis of the latitudinal heat budget, as well as the dependence of the size of the equatorial region on $Ro_{\text{cv,F}}$, is required to form a complete understanding of the $Nu$ scaling in the rotationally-constrained regime. A second, puzzling feature of the rotationally-constrained regime is that the convective velocity scaling (i.e., the $Re$ scaling) is diffusivity-dependent, but the heat transport scaling is not. This discrepancy may simply imply that the r.m.s. fluctuating velocity $U$ (used to compute $Re$) does not accurately characterise the convective velocity at the scale that contributes most to convective heat transport in the rotationally-constrained regime. Alternatively, it may indicate that the dynamics of the CZ are influenced by the heated and cooled regions (by analogy with the influence that boundary layers can have on the bulk in boundary-driven convection; 2000JFM...407...27G; 2012JFM...691..568K; 2016JFM...808..690G). We note that forming a volume-integrated relation between $Re$ and the viscous dissipation rate is one route to obtaining the inertial scaling for $Re$ (Equation 15; 2016JFM...808..690G). As we are interested in the dynamics integrated over the CZ, this integral relation will not be exact, but instead contain ‘boundary terms’ that describe the exchange of kinetic energy between the CZ and the heated and cooled regions. Understanding the details of this exchange may be important for understanding the scaling behaviour of $Re$ in the CZ.
@@
Several aspects of the flows here are unlike those realised in a star like the Sun. We believe most of these discrepancies result from choices and simplifications made in our modeling (rather than from a direct dependence on diffusivity), but testing this assertion will require future work. For example, none of the zonal flows in our calculations are particularly ‘Solar-like’; those that exhibit a fast equator typically also have a ‘polar vortex’, consisting of prograde flow at high latitudes. This may reflect the narrow aspect ratio of the convective shell considered here (see, e.g., 1979ApJ...231..284G). Alternatively, we note previous work has found that the emergence of polar vortices in simulations of shell convection can be dependent on the initial conditions (2014A&A...570A..43K). More generally, we have restricted ourselves to the simplest case of Boussinesq, hydrodynamic convection. This set-up has the advantage of being easily relatable to theory. However, both magnetism and compressibility have important effects on the dynamics of rotating, spherical shell convection (e.g., 2012Icar..219..428G; 2016GeoJI.204.1120Y; 2023A&A...669A..98K; 2015ApJ...798...51H; 2022ApJ...933..199H; 10.1093/mnras/staf1081). Therefore, incorporating these effects is an obvious and important follow-up to this work. In addition, we have not analysed the sensitivity of our results to the functional form of the prescribed internal heating and cooling profile. We intend to study these effects in future work.
@@
Data supporting this manuscript has been uplodaded to Zenodo: zenodo.17961728. We would like to acknowledge an anonymous referee whose feedback helped us to improve this manuscript throughout. We thank Benjamin Brown and Keaton Burns for useful suggestions with respect to configuring and using Dedalus. NTL and TJ-H benefited from stimulating discussions at the workshop Rotating Turbulence: Interplay and Separability of Bulk and Boundary Dynamics, hosted by Institute for Pure and Applied Mathematics (IPAM) at the University of California, Los Angeles. IPAM is supported by the National Science Foundation (grant no. DMS-1925919). MKB thanks the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Frontiers in Dynamo Theory: From the Earth to the Stars (DYT2), where preliminary work on this topic was conducted; this was supported by EPSRC grant no EP/R014604/1. MKB also thanks Evan Anders, and other participants in the DYT2 programme, for useful discussions. NTL, TJ-H, MKB, LKC, and ST received funding from the STFC under grant agreements ST/Y002156/1, ST/Y002296/1, and ST/Y002113/1. LKC also gratefully acknowledges support from the STFC under grant agreement ST/X001083/1, and TJ-H was additionally supported by an STFC Studentship under grant agreement ST/W507453/1. This work used the DiRAC Memory Intensive service (Cosma8) at Durham University, managed by the Institute for Computational Cosmology, and the DiRAC Data Intensive service (DIaL3) at the University of Leicester managed by the University of Leicester Research Computing Service. These facilities are managed on behalf of the STFC DiRAC HPC (www.dirac.ac.uk). The DiRAC services at Durham and Leicester were funded by BEIS, UKRI and STFC capital funding, and STFC operations grants. The service at Durham received funding from Durham University. DiRAC is part of the UKRI Digital Research Infrastructure. {contribution}NTL, MKB, LKC, and ST designed the research. NTL performed the numerical simulations and data analysis. All authors contributed to the interpretation of results and to writing the manuscript.